# Vision Mamba: Efficient Visual Representation Learning with Bidirectional State Space Model

### v2

@@ 3.3 Vim Block
8:     $\mathbf{B}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{B}}_{o}(\mathbf{x}^{\prime}_{o})$
@@
11:     $\mathbf{\Delta}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\log(1+\exp(\mathbf{Linear}^{\mathbf{\Delta}}_{o}(\mathbf{x}^{\prime}_{o})+\mathbf{Parameter}^{\mathbf{\Delta}}_{o}))$
12:     /* shape of $\mathbf{Parameter}^{\mathbf{A}}_{o}$ is $(\mathtt{E},\mathtt{N})$/
13:     $\overline{\mathbf{A}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{Parameter}^{\mathbf{A}}_{o}$
@@
15:     $\mathbf{y}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SSM}(\overline{\mathbf{A}}_{o},\overline{\mathbf{B}}_{o},\mathbf{C}_{o})(\mathbf{x}_{o}^{\prime})$
16:  end for
17:  /* get gated $\mathbf{y}_{o}$/
18:  $\mathbf{y}_{forward}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{forward}\bigodot\mathbf{SiLU}(\mathbf{z})$
19:  $\mathbf{y}_{backward}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{backward}\bigodot\mathbf{SiLU}(\mathbf{z})$
20:  /* residual connection */
21:  $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{Linear}^{\mathbf{T}}(\mathbf{y}_{forward}^{\prime}+\mathbf{y}_{backward}^{\prime})+\mathbf{T}_{l-1}$ Return: $\mathbf{T}_{l}$
Specifically, we present the operations of Vim block in Algo. 21. The input token sequence $\mathbf{T}_{\mathtt{l}-1}$ is first normalized by the normalization layer. Next, we linearly project the normalized sequence to the $\mathbf{x}$ and $\mathbf{z}$ with dimension size $E$. Then, we process the $\mathbf{x}$ from the forward and backward directions. For each direction, we first apply the 1-D convolution to the $\mathbf{x}$ and get the $\mathbf{x}^{\prime}_{o}$. We then linearly project the $\mathbf{x}^{\prime}_{o}$ to the $\mathbf{B}_{o}$, $\mathbf{C}_{o}$, $\mathbf{\Delta}_{o}$, respectively. The $\mathbf{\Delta}_{o}$ is then used to transform the $\overline{\mathbf{A}}_{o}$, $\overline{\mathbf{B}}_{o}$, respectively. Finally, we compute the $\mathbf{y}_{forward}$ and $\mathbf{y}_{backward}$ through the SSM. The $\mathbf{y}_{forward}$ and $\mathbf{y}_{backward}$ are then gated by the $\mathbf{z}$ and added together to get the output token sequence $\mathbf{T}_{\mathtt{l}}$.
3.4 Architecture Details
In summary, the hyper-parameters of our architecture are listed as follows:
[…]
Following ViT [13] and DeiT [60], we first employ 16$\times$16 kernel size projection layer to get a 1-D sequence of non-overlapping patch embeddings. Subsequently, we directly stack $\mathtt{L}$ Vim blocks. By default, we set the number of blocks $\mathtt{L}$ to 24, SSM dimension $\mathtt{N}$ to 16. To align with the model sizes of DeiT series, we set the hidden state dimension $\mathtt{D}$ to 192 and expanded state dimension $\mathtt{E}$ to 384 for the tiny-size variant. For the small-size variant, we set $\mathtt{D}$ to 384 and $\mathtt{E}$ to 768.
@@ 3.5 Efficiency Analysis
Traditional SSM-based methods leverage the fast Fourier transform to boost the convolution operation as shown in Eq. (4). For data-dependent methods, such as Mamba, the SSM operation in Line 11 of Algo. 21 is no longer equivalent to convolution. To address this problem, Mamba and the proposed Vim choose a modern-hardware-friendly way to ensure efficiency. The key idea of this optimization is to avoid the IO-bound and memory-bound of modern hardware accelerators (GPUs).
@@
Computation-Efficiency. SSM in Vim block (Line 11 in Algo.21) and self-attention in Transformer both play a key role in providing global context adaptively. Given a visual sequence $\mathbf{T}\in R^{1\times\mathtt{M}\times\mathtt{D}}$ and the default setting $\mathtt{E}=2\mathtt{D}$, the computation complexity of a global self-attention and SSM are:
@@ 4 Experiment
${}^{\dagger}$ $224^{2}$ 7M 78.3 +2.2
Vim-S $224^{2}$ 26M 80.5
${}^{\dagger}$ $224^{2}$ 26M 81.6 +1.1
Table 1: Comparison with different backbones on ImageNet-1K validation set. ${}^{\dagger}$ represents the model is fine-tuned with our long sequence setting.
@@ 4.1 Image Classification
Settings. We benchmark Vim on the ImageNet-1K dataset [9], which contains 1.28M training images and 50K validation images from 1,000 categories. All models are trained on the training set, and top-1 accuracy on the validation set is reported. For fair comparisons, our training settings mainly follow DeiT [60]. Specifically, we apply random cropping, random horizontal flipping, label-smoothing regularization, mixup, and random erasing as data augmentations. When training on $224^{2}$ input images, we employ AdamW [43] with a momentum of $0.9$, a total batch size of $1024$, and a weight decay of $0.05$ to optimize models. We train the Vim models for $300$ epochs using a cosine schedule, 1×1$\times$1 ×$10^{-3}$ initial learning rate, and EMA. During testing, we apply a center crop on the validation set to crop out $224^{2}$ images. Experiments are performed on 8 A800 GPUs.
@@
Long Sequence Fine-tuning To make full use of the efficient long sequence modeling power of Vim, we continue to fine-tune Vim with a long sequence setting for 30 epochs after ImageNet pretraining. Specifically, we set a patch extraction stride of $8$ while keeping the patch size unchanged, a constant learning rate of $10^{-5}$, and a weight decay of $10^{-8}$.
Results. Tab. 1 compares Vim with ConvNet-based, Transformer-based and SSM-based backbone networks. Compared to ConvNet-based ResNet [24], Vim demonstrates superior performance. For example, when the parameters are roughly similar, the top-1 accuracy of Vim-Small reaches 80.5, which is 4.3 points higher than that of ResNet50. Compared with the conventional self-attention-based ViT [13], Vim outperforms it by considerable margins in terms of both parameter numbers and classification accuracy. When compared to the highly-optimized ViT-variant, i.e., DeiT [60], Vim surpasses it at different scales with comparable parameter numbers: 3.9 points higher for Vim-Tiny over DeiT-Tiny, and 0.7 points higher for Vim-Small over DeiT-Small. Compared with SSM-based S4ND-ViT-B [46], Vim achieves higher top-1 accuracy with 3$\times$ fewer parameters. After long sequence fine-tuning, ${}^{\dagger}$ and ${}^{\dagger}$ all achieve higher results. Among them, ${}^{\dagger}$ even achieves similar results with DeiT-B. The results demonstrate that Vim can be adapted to longer sequence modeling easily and extract stronger visual representation.
@@
[…]
DeiT-Ti
44.4
63.0
47.8
26.1
47.4
61.8
Vim-Ti
45.7
63.9
49.6
26.1
49.0
63.2
Backbone
[…]
DeiT-Ti
38.1
59.9
40.5
18.1
40.5
58.4
Vim-Ti
39.2
60.9
41.7
18.2
41.8
60.2
Table 3: Results of object detection and instance segmentation on the COCO $val$ set using Cascade Mask R-CNN [4] framework.
4.2 Semantic Segmentation
Settings. We conduct experiments for semantic segmentation on the ADE20K [73] and use UperNet [70] as the segmentation framework. We provide detailed settings in Sec. B.
Results. As shown in Tab. 2, Vim consistently outperforms DeiT across different scales: 1.8 mIoU higher for Vim-Ti over DeiT-Ti, and 0.9 mIoU higher for Vim-S over DeiT-S. Compared to the ResNet-101 backbone, our Vim-S achieves the same segmentation performance with nearly 2$\times$ fewer parameters.
Figure 3: GPU memory efficiency comparison between DeiT-Ti [59] and our Vim-Ti on the commonly used downstream framework. We perform batch inference and benchmark the GPU memory on the architecture with the backbone and FPN. Vim requires comparable GPU memory to DeiT with a small resolution, i.e., 512$\times$512. As the input image resolution increases, Vim will use significantly less GPU memory.
4.3 Object Detection and Instance Segmentation
Settings. We conduct experiments for object detection and instance segmentation on the COCO 2017 dataset [38] and use ViTDet [70] as the basic framework. We provide detailed settings in Sec. B.
[…]
We highlight that the accuracy superiority is non-trivial since DeiT is equipped with window attention while Vim works in a pure sequence modeling manner. Specifically, to perform representation learning on high-resolution images (i.e., 1024$\times$1024), we follow ViTDet [37] and modify the DeiT backbone with the use of 2D window attention, which injects 2D prior and breaks the sequential modeling nature of Transformer. Thanks to the efficiency illustrated in Sec. 3.5, Fig. 1 and Fig. 3, we can directly apply Vim on 1024$\times$1024 input images and learn sequential visual representation for object detection and instance segmentation without need for 2D priors in the backbone.
Figure 4: FPS comparison between DeiT-Ti [59] and our Vim-Ti on the commonly used downstream framework. We perform batch inference and benchmark the log-scaled FPS on the architecture with the backbone and FPN. Vim achieves comparable performance to DeiT with a small resolution, i.e., 512$\times$512. As the input image resolution increases, Vim has a higher FPS.
@@ 4.4 Ablation Study
Bidirectional SSM. We ablate the key bidirectional design of Vim, using ImageNet-1K classification and the Segmenter [53] semantic segmentation framework on ADE20K. To fully evaluate the power of learned representation on ImageNet, we use a simple Segmenter head with only 2 layers to perform transfer learning on semantic segmentation. We study these bidirectional strategies:
  • None. We directly adopt the Mamba block to process visual sequence with only the forward direction.
  • Bidirectional Sequence. During training, we randomly flip the visual sequence. This works like data augmentation.
  • Bidirectional Block. We pair the stacked blocks. The first block of each pair processes visual sequence in the forward direction and the second block of each pair processes in the backward direction.
  • Bidirectional SSM. We add an extra SSM for each block to process the visual sequence in the backward direction.
  • Bidirectional SSM + Conv1d. Based on Bidirectional SSM, we further add a backward Conv1d before the backward SSM (Fig. 2).
@@
Classification Design. We ablate the classification design of Vim, benchmarking on ImageNet-1K classification. We study the following classification strategies:
  • Mean pool. We adopt mean pooling on the output feature from the last Vim block and perform classification on this pooled feature.
  • Max pool. We first adapt the classification head on each token of the visual sequence and then perform max pooling on the sequence to get the classification prediction result.
  • Head class token. Following DeiT [60], we concatenate the class token at the head of the visual sequence and perform classification.
  • Double class token. Based on the head class token strategy, we additionally add a class token at the tail of the visual sequence.
  • Middle class token. We add a class token at the middle of the visual sequence and then perform classification on the final middle class token.
@@ Appendix A Visualization
Figure 5: Visualization comparison of DeiT-Ti [60] and our Vim-Ti on the Cascade Mask R-CNN [4] framework. Thanks to the long-range context learning of SSM, we can capture the very large object in the image, which the DeiT-Ti counterpart fails to perceive.
@@ Appendix B Additional Setting
Settings for Semantic Segmentation. We conduct experiments for semantic segmentation on the ADE20K [73] dataset. ADE20K contains 150 fine-grained semantic categories, with 20K, 2K, and 3K images for training, validation, and testing, respectively. We choose UperNet [69] as our base framework. In training, we employ AdamW with a weight decay of $0.01$, and a total batch size of $16$ to optimize models. The employed training schedule uses an initial learning rate of $6$$\times$$10^{-5}$, linear learning rate decay, a linear warmup of $1,500$ iterations, and a total training of $160$K iterations. The data augmentations follow common settings, including random horizontal flipping, random re-scaling within the ratio range $[0.5,2.0]$, and random photometric distortion. During evaluation, we rescale the image to have a shorter side of $512$.
Settings for Object Detection and Instance Segmentation. We conduct experiments for object detection and instance segmentation on the COCO 2017 dataset [38]. The COCO 2017 dataset contains 118K images for training, 5K images for validating, and 20K images for testing. We use the canonical Cascade Mask R-CNN [4] as the base framework. For ViT-based backbones, we apply extra configurations (e.g., interleaved window & global attention) to handle the high-resolution images following ViTDet [37]. For SSM-based Vim, we directly use it without any modifications. Other training and evaluation settings are just the same. During training, we employ AdamW with a weight decay of $0.1$, and a total batch size of $64$ to optimize models. The employed training schedule uses an initial learning rate of $1$$\times$$10^{-4}$, linear learning rate decay, and a total training of $380$K iterations. The data augmentations use large-scale jitter data augmentation [18] to 1024$\times$1024 input images. During evaluation, we rescale the image to have a shorter side of 1024.

### v3

@@ 3.4 Architecture Details
8:    $\mathbf{B}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}_{o}^{\mathbf{B}}(\mathbf{x}^{\prime}_{o})$
@@
11:    $\mathbf{\Delta}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\log(1+\exp(\mathbf{Linear}_{o}^{\mathbf{\Delta}}(\mathbf{x}^{\prime}_{o})+\mathbf{Parameter}_{o}^{\mathbf{\Delta}}))$
12:    /* shape of $\mathbf{Parameter}_{o}^{\mathbf{A}}$ is $(\mathtt{E},\mathtt{N})$/
13:    $\overline{\mathbf{A}_{o}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{Parameter}_{o}^{\mathbf{A}}$
@@
15:    /* initialize $h_{o}$ and $\mathbf{y}_{o}$ with 00 */
16:    $h_{o}$ : $(\mathtt{B},\mathtt{E},\mathtt{N})$leftarrow zeros $(\mathtt{B},\mathtt{E},\mathtt{N})$
17:    $\mathbf{y}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow zeros $(\mathtt{B},\mathtt{M},\mathtt{E})$
18:    /* SSM recurrent */
19:    for $i$ in {0, …, M-1} do
20:       $h_{o}$ = $\overline{\mathbf{A}_{o}}[:,i,:,:]\bigodot h_{o}+\overline{\mathbf{B}_{o}}[:,i,:,:]\bigodot\mathbf{x}_{o}^{\prime}[:,i,:,{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathtt{None}}]$
21:       $\mathbf{y}_{o}[:,i,:]$ = $h_{o}\bigotimes\mathbf{C}_{o}[:,i,:]$
22:    end for
23:  end for
24:  /* get gated $\mathbf{y}$/
25:  $\mathbf{y}_{forward}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{forward}\bigodot\mathbf{SiLU}(\mathbf{z})$
26:  $\mathbf{y}_{backward}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{backward}\bigodot\mathbf{SiLU}(\mathbf{z})$
27:  /* residual connection */
28:  $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{Linear}^{\mathbf{T}}(\mathbf{y}_{forward}^{\prime}+\mathbf{y}_{backward}^{\prime})+\mathbf{T}_{l-1}$
29:  Return: $\mathbf{T}_{l}$
@@ 3.5 Efficiency Analysis
Traditional SSM-based methods leverage the fast Fourier transform to boost the convolution operation as shown in Eq. (3). For data-dependent methods, such as Mamba, the SSM operation in Line 11 of Algo. 29 is no longer equivalent to convolution. To address this problem, Mamba and the proposed Vim choose a modern-hardware-friendly way to ensure efficiency. The key idea of this optimization is to avoid the IO-bound and memory-bound of modern hardware accelerators (GPUs).
@@
Computation-Efficiency. SSM in Vim block (Line 11 in Algo.29) and self-attention in Transformer both play a key role in providing global context adaptively. Given a visual sequence $\mathbf{T}\in R^{1\times\mathtt{M}\times\mathtt{D}}$ and the default setting $\mathtt{E}=2\mathtt{D}$, the computation complexity of a global self-attention and SSM are:
@@ 4 Experiment
Vim-Ti $224^{2}$ 7M 78.3 +2.2
Vim-S $224^{2}$ 26M 80.3
Vim-S $224^{2}$ 26M 81.4 +1.1
Vim-B $224^{2}$ 98M 81.9
Vim-B $224^{2}$ 98M 83.2 +1.3
Table 1: Comparison with different backbones on ImageNet-1K validation set. represents the model is fine-tuned with our long sequence setting.
@@ 4.1 Image Classification
Settings. We benchmark Vim on the ImageNet-1K dataset (Deng et al., 2009), which contains 1.28M training images and 50K validation images from 1,000 categories. All models are trained on the training set, and top-1 accuracy on the validation set is reported. For fair comparisons, our training settings mainly follow DeiT (Touvron et al., 2021b). Specifically, we apply random cropping, random horizontal flipping, label-smoothing regularization, mixup, and random erasing as data augmentations. When training on $224^{2}$ input images, we employ AdamW (Loshchilov & Hutter, 2019) with a momentum of $0.9$, a total batch size of $1024$, and a weight decay of $0.05$ to optimize models. We train the Vim models for $300$ epochs using a cosine schedule, 1×1$\times$1 ×$10^{-3}$ initial learning rate, and EMA. During testing, we apply a center crop on the validation set to crop out $224^{2}$ images. Experiments are performed on 8 A800 GPUs.
@@
Long Sequence Fine-tuning To make full use of the efficient long sequence modeling power of Vim, we continue to fine-tune Vim with a long sequence setting for 30 epochs after pretraining. Specifically, we set a patch extraction stride of $8$ while keeping the patch size unchanged, a constant learning rate of $10^{-5}$, and a weight decay of $10^{-8}$.
Results. Tab. 1 compares Vim with ConvNet-based, Transformer-based and SSM-based backbone networks. Compared to ConvNet-based ResNet (He et al., 2016), Vim demonstrates superior performance. For example, when the parameters are roughly similar, the top-1 accuracy of Vim-Small reaches 80.3, which is 4.1 points higher than that of ResNet50. Compared with the conventional self-attention-based ViT (Dosovitskiy et al., 2020), Vim outperforms it by considerable margins in terms of both parameter numbers and classification accuracy. When compared to the highly-optimized ViT-variant, i.e., DeiT (Touvron et al., 2021b), Vim surpasses it at different scales with comparable parameter numbers: 3.9 points higher for Vim-Tiny over DeiT-Tiny, 0.5 points higher for Vim-Small over DeiT-Small, and 0.1 points higher for Vim-Base over DeiT-Base. Compared with SSM-based S4ND-ViT-B (Nguyen et al., 2022), Vim achieves similar top-1 accuracy with 3$\times$ fewer parameters. After long sequence fine-tuning, Vim-Tiny, Vim-S, and Vim-B all achieve higher results. Among them, Vim-S even achieves similar results with DeiT-B. The results demonstrate that Vim can be adapted to longer sequence modeling easily and extract stronger visual representation.
@@
DeiT-Ti 38.1 59.9 40.5 18.1 40.5 58.4
Vim-Ti 39.2 60.9 41.7 18.2 41.8 60.2
Table 3: Results of object detection and instance segmentation on the COCO $val$ set using Cascade Mask R-CNN (Cai & Vasconcelos, 2019) framework.
Figure 3: FPS comparison between DeiT-Ti (Touvron et al., 2021a) and our Vim-Ti on the commonly used downstream framework. We perform batch inference and benchmark the log-scaled FPS on the architecture with the backbone and FPN. Vim achieves comparable performance to DeiT with a small resolution, i.e., 512$\times$512. As the input image resolution increases, Vim has a higher FPS.
4.2 Semantic Segmentation
Settings. We conduct experiments for semantic segmentation on the ADE20K (Zhou et al., 2019) and use UperNet (Xiao et al., 2018b) as the segmentation framework. We provide detailed settings in Sec. B. Results. As shown in Tab. 2, Vim consistently outperforms DeiT across different scales: 1.8 mIoU higher for Vim-Ti over DeiT-Ti, and 0.9 mIoU higher for Vim-S over DeiT-S. Compared to the ResNet-101 backbone, our Vim-S achieves the same segmentation performance with nearly 2$\times$ fewer parameters.
4.3 Object Detection and Instance Segmentation
Settings. We conduct experiments for object detection and instance segmentation on the COCO 2017 dataset (Lin et al., 2014) and use ViTDet (Xiao et al., 2018b) as the basic framework. We provide detailed settings in Sec. B.
Figure 4: GPU memory efficiency comparison between DeiT-Ti (Touvron et al., 2021a) and our Vim-Ti on the commonly used downstream framework. We perform batch inference and benchmark the GPU memory on the architecture with the backbone and FPN. Vim requires comparable GPU memory to DeiT with a small resolution, i.e., 512$\times$512. As the input image resolution increases, Vim will use significantly less GPU memory.
@@ 4.3 Object Detection and Instance Segmentation
We highlight that the accuracy superiority is non-trivial since DeiT is equipped with window attention while Vim works in a pure sequence modeling manner. Specifically, to perform representation learning on high-resolution images (i.e., 1024$\times$1024), we follow ViTDet (Li et al., 2022c) and modify the DeiT backbone with the use of 2D window attention, which injects 2D prior and breaks the sequential modeling nature of Transformer. Thanks to the efficiency illustrated in Sec. 3.5, Fig. 1 and Fig. 4, we can directly apply Vim on 1024$\times$1024 input images and learn sequential visual representation for object detection and instance segmentation without need for 2D priors in the backbone.
@@ 4.4 Ablation Study
Bidirectional SSM. We ablate the key bidirectional design of Vim, using ImageNet-1K classification and the Segmenter (Strudel et al., 2021) semantic segmentation framework on ADE20K. To fully evaluate the power of learned representation on ImageNet, we use a simple Segmenter head with only 2 layers to perform transfer learning on semantic segmentation. We study the following bidirectional strategies. None: We directly adopt the Mamba block to process visual sequence with only the forward direction. Bidirectional Sequence: During training, we randomly flip the visual sequence. This works like data augmentation. Bidirectional Block: We pair the stacked blocks. The first block of each pair processes visual sequence in the forward direction and the second block of each pair processes in the backward direction. Bidirectional SSM: We add an extra SSM for each block to process the visual sequence in the backward direction. Bidirectional SSM + Conv1d: Based on Bidirectional SSM, we further add a backward Conv1d before the backward SSM (Fig. 2).
@@
Classification Design. We ablate the classification design of Vim, benchmarking on ImageNet-1K classification. We study the following classification strategies. Mean pool: We adopt mean pooling on the output feature from the last Vim block and perform classification on this pooled feature. Max pool: We first adapt the classification head on each token of the visual sequence and then perform max pooling on the sequence to get the classification prediction result. Head class token: Following DeiT (Touvron et al., 2021b), we concatenate the class token at the head of the visual sequence and perform classification. Double class token: Based on the head class token strategy, we additionally add a class token at the tail of the visual sequence. Middle class token: We add a class token at the middle of the visual sequence and then perform classification on the final middle class token.
@@ 5 Conclusion and Future Work
Impact Statement
We advance the efficiency of the generic vision backbone. Any societal consequences or impacts that typically relate to work focused on increased efficiency also apply here, as such work necessarily improves the practicality of vision backbone for an array of visual applications with high-resolution input images.
@@ Acknowledgement
This work was partially supported by the National Science and Technology Major Project under Grant No. 2023YFF0905400 and National Natural Science Foundation of China (NSFC) under Grant No. 62276108.
@@ Appendix A Visualization
Figure 5: Visualization comparison of DeiT-Ti (Touvron et al., 2021b) and our Vim-Ti on the Cascade Mask R-CNN (Cai & Vasconcelos, 2019) framework. Thanks to the long-range context learning of SSM, we can capture the very large object in the image, which the DeiT-Ti counterpart fails to perceive.
@@ Appendix B Additional Setting
Settings for Semantic Segmentation. We conduct experiments for semantic segmentation on the ADE20K (Zhou et al., 2019) dataset. ADE20K contains 150 fine-grained semantic categories, with 20K, 2K, and 3K images for training, validation, and testing, respectively. We choose UperNet (Xiao et al., 2018a) as our base framework. In training, we employ AdamW with a weight decay of $0.01$, and a total batch size of $16$ to optimize models. The employed training schedule uses an initial learning rate of $6$$\times$$10^{-5}$, linear learning rate decay, a linear warmup of $1,500$ iterations, and a total training of $160$K iterations. The data augmentations follow common settings, including random horizontal flipping, random re-scaling within the ratio range $[0.5,2.0]$, and random photometric distortion. During evaluation, we rescale the image to have a shorter side of $512$.
Settings for Object Detection and Instance Segmentation. We conduct experiments for object detection and instance segmentation on the COCO 2017 dataset (Lin et al., 2014). The COCO 2017 dataset contains 118K images for training, 5K images for validating, and 20K images for testing. We use the canonical Cascade Mask R-CNN (Cai & Vasconcelos, 2019) as the base framework. For ViT-based backbones, we apply extra configurations (e.g., interleaved window & global attention) to handle the high-resolution images following ViTDet (Li et al., 2022c). For SSM-based Vim, we directly use it without any modifications. Other training and evaluation settings are just the same. During training, we employ AdamW with a weight decay of $0.1$, and a total batch size of $64$ to optimize models. The employed training schedule uses an initial learning rate of $1$$\times$$10^{-4}$, linear learning rate decay, and a total training of $380$K iterations. The data augmentations use large-scale jitter data augmentation (Ghiasi et al., 2021) to 1024$\times$1024 input images. During evaluation, we rescale the image to have a shorter side of 1024.
Appendix C Extended Comparison on Hierarchical Architecture
To further compare with hierarchical architectures, we propose another variant Hier-Vim by replacing shifted local window attention in SwinTransformer with the proposed global bidirectional SSM. We detail the configuration in Tab. 6
Model #Blocks #Channels Params
Hier-Vim-T [2, 2, 5, 2] [96, 192, 384, 768] 30M
Hier-Vim-S [2, 2, 15, 2] [96, 192, 384, 768] 50M
Hier-Vim-B [2, 2, 15, 2] [128, 256, 512, 1024] 89M
Table 6: Detailed configurations of different variants of Hier-Vim. We provide the number of channels and blocks in 4 stages.
Method
image
size
#param.
ImageNet
top-1 acc.
Swin-T (Liu et al., 2021) $224^{2}$ 28M 81.2
FocalTransformer-T (Yang et al., 2021) $224^{2}$ 29M 82.2
CVT-21 (Wu et al., 2021) $224^{2}$ 32M 82.5
MetaFormer-S35 (Yu et al., 2022) $224^{2}$ 31M 81.4
GFNet-H-S (Rao et al., 2021) $224^{2}$ 32M 81.5
Hier-Vim-T $224^{2}$ 30M 82.5
Swin-S (Liu et al., 2021) $224^{2}$ 50M 83.2
FocalTransformer-S (Yang et al., 2021) $224^{2}$ 51M 83.5
MetaFormer-S35 (Yu et al., 2022) $224^{2}$ 73M 82.5
GFNet-H-B (Rao et al., 2021) $224^{2}$ 54M 82.9
Hier-Vim-S $224^{2}$ 50M 83.4
Swin-B (Liu et al., 2021) $224^{2}$ 88M 83.5
FocalTransformer-B (Yang et al., 2021) $224^{2}$ 90M 83.8
Hier-Vim-B $224^{2}$ 89M 83.9
Table 7: Comparison with hierarchical architectures on ImageNet-1K validation set.
Classification on ImageNet. Following the standard training and validation protocols (Liu et al., 2021, 2024), we compare Hier-Vim with popular hierarchical architectures across tiny, small, and base model sizes in Tab. 7. The results indicate that Hier-Vim outperforms Swin Transformer by 1.3% at the tiny size, 0.2% at the small size, and 0.4% at the base size, demonstrating competitive performance against well-established and highly-optimized modern hierarchical architectures.